I need help on how to solve this answer! I just want to see the work so I know what to do from now on. Or maybe you can just explain how to solve these
y = 2x+ 5
2x - 5y = 4

I need to solve this with the elimination method. please help

Answer by ankor@dixie-net.com(22740)

y = 2x + 5
2x - 5y = 4
we need both equations in the standard form, rearrange the 1st equation
then add the two equations
-2x + y = 5
2x - 5y = 4
------------------addition eliminates x, find y
-4y = 9
y = 9/-4
y = -2.25
:
Find y using the 1st original equation; y = 2x + 5
replace y with -2.25
-2.25 = 2x + 5
-2.25 - 5 = 2x
-7.25 = 2x
x = -7.25/2
x = -3.265
:
:
Check our solutions in the 2nd equation 2x - 5y = 4
replace x and y
2(-3.265) - 5(-2.25) = 4
-7.25 + 11.25 = 4; confirms our solutions of x=-3.265; y=-2.25